Find the area of the following circles, given that:
 Radius = 5 cm

Grade:12

1 Answers

Harshit Singh
2 years ago
Dear Student

Given, radius of circle = 5 cm
Then,
Area of the circle =πr^2
= 22/7 × 5^2
= 22/7 × 25
= 550/7
= 78.57 cm^2

Thanks